### v2

# Towards Optimal Aggregation of Varying Range Dependencies in Haze Removal

Abstract
Haze removal aims to restore a clear image from a hazy input. Existing methods have shown significant efficacy by capturing either short-range dependencies for local detail preservation or long-range dependencies for global context modeling. Given the complementary strengths of both approaches, a intuitive advancement is to explicitly integrate them into a unified framework. However, this potential remains underexplored in current research. In this paper, we propose DehazeMatic, which leverages the proposed Transformer-Mamba Dual Aggregation block to simultaneously and explicitly captures both short- and long-range dependencies through dual-path design for improved restoration. To ensure that dependencies at varying ranges contribute optimally to performance, we conduct extensive experiments to identify key influencing factors and determine that an effective aggregation mechanism should be guided by the joint consideration of haze density and semantic information. Building on these insights, we introduce the CLIP-enhanced Dual-path Aggregator, which utilizes the rich semantic priors encapsulated in CLIP and the estimated haze density map, derived from its powerful generalization ability, to instruct the aggregation process. Extensive experiments demonstrate that DehazeMatic outperforms sort-of-the-art methods across various benchmarks.
Figure 1: Illustration of factors influencing the relative importance between short-range and long-range dependencies. We train a Transformer-based dehazing network with a global receptive field and synthesis hazy images with different densities based on the Atmospheric Scattering Model [49]. (a) The horizontal axis represents the Euclidean distance between the current token and all other tokens, while the vertical axis represents the average attention weights between the current token and all tokens at the corresponding distance. The curve data is derived from all tokens of all images in the test set. (b) The bottom visualizes the attention map between the chosen tokens (indicated by the tails of red arrows) and all other tokens. On the left is the visualization of tokens at the same position corresponding to the same ground truth image, but with different haze densities. On the right is a visualization of tokens with different semantic information on the same homogeneous haze image. More analysis is in supplementary material.
1 Introduction
Single image dehazing aims to recover a clear image from a hazy input, serving as a crucial pre-processing step for high-level vision tasks under hazy conditions, such as object detection [39] and semantic segmentation [53].
Figure 2: Illustration of CLIP [52]’s potential to perceive haze and its density. We use the ViT-B/32 OpenCLIP [26] model, pre-trained on the LAION dataset, for our experiments. The values above the images indicate the PSNR and SSIM metrics, which reflect haze density from a quantitative perspective. The values below the images represent the similarity scores between the images and the paired prompts from CLIP. Notably, as haze density increases, the similarity score with the haze-describing prompt also increases.
Current data-driven methods can be categorized into two classes based on the receptive field size of the operators used for feature extraction, namely (i) using convolution [5, 7, 13, 37, 55, 56, 72] or window-based self-attention [35, 34, 59, 64] and (ii) using linear self-attention [51] or State Space Model (SSM) [58, 80, 73]. The former excels at capturing short-range dependencies but lacks effective global modeling capabilities [30, 62, 12]. In contrast, the latter directly models long-range dependencies while ensuring efficient computation; however, it fails to preserve local 2D inductive bias [25, 14].
Both categories of methods have achieved excellent performance, and their respective strengths can address the limitations of the other. Therefore, a natural idea is that explicitly combining short- and long-range dependencies (i.e., leveraging the strengths of both approaches) is expected to yield even more promising results. However, this integration has not been well explored. To address this gap, we will employ a dual-path design that concurrently captures fine-grained perception from nearest neighbor features and holistic contextual cues from non-local relationships. The ensuing problem is that reasonably integrating short- and long-range dependencies from these two paths is a non-trivial task, because tokens with different characteristics in the image require varying ranges of dependencies. Simple addition or concatenation often leads to suboptimal performance. More importantly, there is still no clear guidance on how to determine the relative importance between short-range and long-range dependencies for each token.
One preliminary hypothesis is that haze density serves as a determining factor. To verify this, we conduct an experiment. As shown in Figure 1(a), higher haze density increases reliance on long-range dependencies, whereas lower haze density requires more short-range dependencies. Furthermore, additional investigation reveals that tokens with varying semantic information require different ranges of dependencies, as illustrated on the right side of Figure 1(b). Based on these findings, we conclude that haze density and semantic information jointly determine the relative importance of short-range and long-range dependencies.
In this paper, building on the above insights, we propose DehazeMatic, where Mamba and Transformer are Integrated by CLIP for image dehazing. The core building block of the DehazeMatic is the proposed Transformer-Mamba Dual Aggregation (TrambaDA) block, which employs a dual-path design. In one path, we utilize window-based self-attention [45] to capture short-range dependencies. In the other path, inspired by recent advances in SSMs [18, 19, 44, 82, 42], we incorporate the Selective Scan Space State Sequential Model (S6) block [18], which maintains global information indefinitely with linear complexity. We also adopt the four-way scanning strategy to mitigate the “direction-sensitive" problem caused by the non-causal nature of visual data [44].
Next, to ensure that both short- and long-range dependencies contribute optimally to performance, a dual-path aggregator capable of adaptively learning the relative importance of each path is required. As indicated by Figure 1, the aggregator must capture both the haze density and semantic information of each token in the input image. Recent advances in CLIP [52], pre-trained on web-scale datasets, reveal that it not only encapsulates rich semantic priors (e.g., enabling strong performance in zero-shot semantic segmentation [81, 76]), but also has the potential to perceive haze and its density (as shown in Figure 2). Motivated by this, we propose the CLIP-enhanced Dual-path Aggregator (CedA). Specifically, we treat the affinity scores between the the patch-wise image embeddings from the CLIP image encoder and the text embeddings of paired prompts describing clear and hazy conditions from the CLIP text encoder as the estimated haze density map. To find the best prompts for better estimation and avoid labor-intensive prompt engineering, these paired prompts are learned. For semantic information, we reuse the the patch-wise image embeddings to provide high-level and abstract features while avoiding additional computational overhead. To supplement the low-level details, we incorporate the input features from the TrambaDA blocks and align both representations using a bidirectional cross-attention mechanism. Finally, we generate dual-path aggregation weights based on the refined semantic information and the estimated haze density map.
Figure 3: (a) The architecture of our proposed DehazeMatic. (b) The inner structure of Transformer-Mamba Dual Aggregation (TrambaDA) block. The weights generated by CLIP-enhanced Dual-path Aggregator (CedA) are used to aggregate the outputs of dual-path.
Our contributions can be summarized as follows:
  • We propose DehazeMatic, which demonstrates the superiority of explicitly and simultaneously capturing both short- and long-range dependencies while effectively aggregating dependencies across varying ranges. It achieves SOTA performance in various benchmarks.
  • For the first time, we identify the key factors required for reasonably aggregating dependencies across different ranges in image dehazing. Based on these insights, we propose the CedA module, which performs dynamic aggregation according to haze density and semantic characteristics of each token within the image.
  • The proposed CedA module further explore the potential of CLIP in image dehazing and, for the first time, demonstrate its capability in haze density estimation, all without the need for fine-tuning the encoder.
2 Related Work
2.1 Single Image Dehazing
Due to spatially variant transmission map and atmospheric light, single image dehazing is a highly ill-posed problem. Early prior-based methods [23, 16, 31, 60, 83, 6] achieved dehazing by imposing various assumptions to estimate key parameters in the Atmospheric Scattering Model (ASM) [49]. ASM provides strong theoretical support and interpretability for these methods; however, they often fail when images deviate from statistical laws.
With the rapid advancement of deep learning [33], numerous learning-based methods have been proposed, resulting in improved performance. Early methods [7, 37] employ neural networks to estimate key parameters in the ASM and subsequently restore the haze-free images. Later, ASM-independent deep networks [54, 55, 43, 40, 57, 13, 75, 50, 36, 66, 70, 59, 17, 68, 77, 8, 15, 9, 65, 69, 11] directly estimate clear images or haze residuals. DehazeFormer [59] is a representative method that achieves efficient feature extraction through window-based self-attention and several key modifications. However, its inherently limited receptive field constrains its performance potential. To achieve a global receptive field with lower computational overhead, some methods [58, 71] incorporate frequency domain features into image dehazing, offering the additional benefit of more direct learning of high-frequency information. For the same purpose, other researchers [79, 73] introduce the Mamba [18] architecture into image dehazing. However, none of these methods address how to better integrate global semantic features with local texture features.
2.2 CLIP for Low-Level Vision Tasks
Classic vision-language models, such as CLIP [52], aim to learn aligned features in the embedding space from image-text pairs using contrastive learning. Some studies have explored leveraging the rich prior knowledge encapsulated in CLIP to assist with low-level vision tasks.
In All-in-One image restoration, some researchers [47, 1, 27] use degradation embeddings from the CLIP image encoder to implicitly guide the network in making adaptive responses. In monocular depth estimation, recent studies [74, 4, 24] use CLIP to map patches of the input image to specific semantic distance tokens, which are then projected onto a quantified depth bin for estimation. In low-light enhancement, some methods [41, 48] use text-image similarity between the enhanced results and learnable prompt pairs to train the enhancement network.
Figure 4: Illustration of the CLIP-enhanced Dual-path Aggregator generating aggregation weights. To ensure computational efficiency, the TrambaDA blocks share the same aggregation weights at each stage.
Some works also introduce CLIP into image [63] and video [53] dehazing, but they all use text-image similarity between dehazed results and contrastive prompt sets as regularization to guide the restoration process. In contrast, our method further exploits the potential of CLIP by directly incorporating its latent embeddings into the main network to guide the dehazing process.
3 Method
3.1 Overall Architecture
The architecture of DehazeMatic is depicted in Figure 3. Given a hazy image, shallow features are first extracted and then passed through a symmetric encoder-decoder framework. Each stage of encoder and decoder consists of multiple Transformer-Mamba Dual Aggregation (TrambaDA) blocks, along with corresponding downsampling or upsampling operations. We also incorporate skip connections between the encoder and decoder at each level. Finally, the output of the last decoder stage is combined with the input hazy image to produce the haze-free image.
3.2 Transformer-Mamba Dual Aggregation Block
The TrambaDA block’s backbone employs a dual-path design, with one path capturing short-range dependencies and the other capturing long-range dependencies, thereby providing complementary restoration information for image tokens with distinct characteristics.
Capturing Short-range Dependencies.  We construct this path using window-based self-attention [45], which offers stronger fitting capability through dynamic weights. Specifically, given the input feature $F_{\rm{in}}\in\mathbb{R}^{H\times W\times C}$, we first divide $F_{\rm{in}}$ into non-overlapping windows with size of $M\times M$, then obtain flattened feature $F_{\mathrm{in}}^{i}\in\mathbb{R}^{M^{2}\times C}$ of each window $i$. For simplicity, we only consider the single-head attention. Performing self-attention on non-overlapping windows can be formulated as:
$\begin{split}F_{\mathrm{out\_{short}}}^{i}=&\text{Attention}(F_{\mathrm{in}}^{i}W^{Q},F_{\mathrm{in}}^{i}W^{K},F_{\mathrm{in}}^{i}W^{V}),\\ \text{Attention}&(Q,K,V)=\text{Softmax}(\frac{{QK}^{T}}{\sqrt{d_{k}}})V,\\ F_{\mathrm{out\_{short}}}=&\{F_{\mathrm{out\_{short}}}^{1},F_{\mathrm{out\_{short}}}^{2},\cdots,F_{\mathrm{out\_{short}}}^{N}\},\end{split}$ (1)
where $N=HW/M^{2}$; $W^{Q},W^{K},W^{V}\in\mathbb{R}^{C\times d_{k}}$ are projection matrices for queries, keys, and values, respectively.
Figure 5: Illustration of estimated haze density maps. The proposed method can identify the relative haze density not only in different regions of non-homogeneous haze images but also between two homogeneous haze images.
Capturing Long-range Dependencies.  We build this path using Mamba [18]’s S6 block, which preserves the global receptive field with linear complexity and incorporates hardware-aware algorithms.
Due to the “direction-sensitive” problem caused by the non-causal nature of visual data, directly applying the S6 block on the flattened image will result in suboptimal performance [44]. Following Vmamba [44], we unfold feature map along four directions, generating four independent sequences. Each feature sequence is processed through the S6 block, and all output sequences are merged into a new sequence. This process can be represented as follows:
$\begin{split}F_{\mathrm{in}}^{d}=expand&(F_{\mathrm{in}},d),d=\{1,2,3,4\},\\ \bar{F}&{}_{\mathrm{in}}^{d}=\mathrm{S6}(F_{\mathrm{in}}^{d}),\\ F_{\mathrm{out\_{long}}}=m&erge(\bar{F}_{\mathrm{in}}^{1},\bar{F}_{\mathrm{in}}^{2},\bar{F}_{\mathrm{in}}^{3},\bar{F}_{\mathrm{in}}^{4}),\end{split}$ (2)
where $expand(\cdot)$ and $merge(\cdot)$ correspond to the scan expand and scan merge operations, respectively.
3.3 CLIP-enhanced Dual-path Aggregator
The CLIP-enhanced Dual-path Aggregator (CedA) utilizes the generated pixel-level weights to adaptively aggregate dual-path outputs, ensuring that dependencies across varying ranges contribute optimally to performance:
$\begin{split}F_{out}=W_{short}F_{\mathrm{out\_{short}}}+W_{long}F_{\mathrm{out\_{long}}}.\end{split}$ (3)
$W_{short},W_{long}\in\mathbb{R}^{H\times W}$ are the aggregation weights corresponding to short- and long-range dependencies, respectively, and their generation process is shown in Figure 4.
3.3.1 Estimating Haze Density Map
Define Estimation Pipeline.  Inspired by Figure 2, we leverage CLIP to estimate the haze density map. Specifically, given a hazy image $I_{\mathrm{haze}}\in\mathbb{R}^{H\times W\times 3}$ and a paired prompts describing haze and clear conditions $T=[T_{\mathrm{haze}},T_{\mathrm{clear}}]$ (e.g., [’hazy image’, ’clear image’]), we encode them into latent space using a pretrained CLIP model:
$\begin{split}F_{\mathrm{image}}=&\Phi_{\mathrm{image}}(I_{\mathrm{haze}})\in\mathbb{R}^{H_{p}\times W_{p}\times d_{c}},\\ &F_{\mathrm{text}}=\Phi_{\mathrm{text}}(T)\in\mathbb{R}^{2\times d_{c}},\end{split}$ (4)
where $\Phi_{\mathrm{text}}$ is the CLIP text encoder, and $\Phi_{\mathrm{image}}$ is the CLIP image encoder without the final pooling layer, allowing the generation of a patch-wise density map. $H_{p}$ and $W_{p}$ denote the height and width of the encoded image embedding, respectively, while $d_{c}$ denotes the hidden dimension of CLIP. As shown in Figure 2, as the haze density increases, the similarity scores between $F_{\mathrm{image}}$ and the haze-describing prompt $T_{\mathrm{haze}}$ also increase. Therefore, we treat these similarity scores as the estimated density map $\mathcal{M}_{d}$:
$\begin{split}\mathcal{M}_{d}=\mathrm{Softmax}(\mathrm{sim}(F_{\mathrm{image}},F_{\mathrm{text}}))[:,:,0],\end{split}$ (5)
where $\mathrm{sim}(\cdot,\cdot)$ is the similarity calculation, and final haze density map $\mathcal{M}_{d}\in\mathbb{R}^{H_{p}\times W_{p}}$.
Table 1: Quantitative comparisons of various methods on various dehazing benchmarks. The best results and the second best results are in bold and underline, respectively.
Methods Publication RESIDE SOTS (Homogeneous Haze) Non-homogeneous Haze Overhead
Outdoor Indoor Average NH-Haze Dense-Haze
$\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$Param MACs
DCP [22] TPAMI-10 19.14 0.861 16.61 0.855 17.88 0.858 10.57 0.520 10.06 0.385 - -
AOD-Net [37] ICCV-17 24.14 0.920 20.51 0.816 22.33 0.868 15.40 0.569 13.14 0.414 1.76K 0.12G
GridDehazeNet [43] ICCV-19 30.86 0.982 32.16 0.984 31.51 0.983 18.33 0.667 14.96 0.533 0.96M 21.55G
FFA-Net [50] AAAI-20 33.57 0.984 36.39 0.989 34.98 0.987 19.87 0.692 16.09 0.503 4.46M 288.86G
DeHamer [20] CVPR-22 35.18 0.986 36.63 0.988 35.91 0.987 20.66 0.684 16.62 0.560 132.45M 59.25G
SRDefog [28] ACCV-22 - - - - - - 20.99 0.610 16.67 0.500 12.56M 24.18M
MAXIM-2S [61] CVPR-22 34.19 0.985 38.11 0.991 36.15 0.988 - - - - 14.10M 216.00G
SGID-PFF [5] TIP-22 30.20 0.975 38.52 0.991 34.36 0.983 - - - - 13.87M 156.67G
PMNet [70] ECCV-22 34.74 0.985 38.41 0.990 36.58 0.988 20.42 0.730 16.79 0.510 18.90M 81.13G
MB-TaylorFormer-B [51] ICCV-23 37.42 0.989 40.71 0.992 39.07 0.991 - - 16.66 0.560 2.68M 38.50G
MITNet [58] MM-23 35.18 0.988 40.23 0.992 37.71 0.990 21.26 0.712 16.97 0.606 2.73M 16.42G
DehazeFormer [59] TIP-23 34.29 0.983 38.46 0.994 36.38 0.989 20.31 0.761 16.66 0.595 4.63M 48.64G
SCANet [21] CVPR-23 - - - - - - 19.52 0.649 15.35 0.508 2.39M 258.63G
DEANet [8] TIP-24 36.03 0.989 40.20 0.993 38.12 0.991 20.84 0.801 16.73 0.602 3.65M 32.23G
UVM-Net [79] Arxiv-24 34.92 0.984 40.17 0.996 37.55 0.990 - - - - 19.25M 173.55G
OKNet [10] AAAI-24 35.45 0.992 37.59 0.994 36.52 0.993 20.29 0.800 16.85 0.620 4.42M 39.54G
DCMPNet [77] CVPR-24 36.56 0.993 42.18 0.996 39.37 0.995 - - - - 18.59M 80.42G
DehazeMatic - 38.21 0.995 41.50 0.996 39.86 0.996 21.47 0.806 17.28 0.629 4.58M 35.50G
Train Learnable Haze/Clear Prompts.  To find the best prompts for improved estimation and avoid time-consuming prompt engineering, we use learnable prompt tokens instead of manually predefined prompts to describe abstract haze and clear conditions.
The training process consists of two stages. In the first stage, we optimize the learnable paired prompts using cross-entropy loss, enabling them to preliminarily distinguish between haze and clear conditions. Given a hazy and clear image $I_{\mathrm{haze}},I_{\mathrm{clear}}\in\mathbb{R}^{H\times W\times 3}$, we define the first-stage loss $\mathcal{L}_{1}$ as:
$\begin{split}\mathcal{L}_{1}=-(y*\log(\hat{y})+(1-y)*\log(1-\hat{y})),\\ \hat{y}=\frac{e^{cos(\mathrm{Pool}(\Phi_{\mathrm{image}}(I)),\Phi_{\mathrm{text}}(T_{\mathrm{clear}}))}}{\sum_{i\in\{\mathrm{haze},\mathrm{clear}\}}e^{cos(\mathrm{Pool}(\Phi_{\mathrm{image}}(I)),\Phi_{\mathrm{text}}(T_{i}))}},\end{split}$ (6)
where $I\in\{I_{\mathrm{haze}},I_{\mathrm{clear}}\}$ and $y$ is the label of the current image, 0 is for hazy image $I_{\mathrm{haze}}$ and 1 is for clear image $I_{\mathrm{clear}}$. Since $\Phi_{\mathrm{image}}$ is the CLIP image encoder without the final pooling layer, we apply a global pooling layer $\mathrm{Pool}()$ to obtain the latent image embedding with a shape of $\mathbb{R}^{1\times d_{c}}$.
In the second stage, our goal is to achieve better prediction of haze density. The most straightforward and effective optimization approach is regression. However, there is no ground-truth density map corresponding to the hazy image in the existing dataset. To address this, we construct triplet data $\{I_{\mathrm{haze}},I_{\mathrm{clear}},I_{\mathrm{density}}\}$ based on Atmospheric Scattering Model [49]. The second-stage loss $\mathcal{L}_{2}$ is defined as:
$\mathcal{L}_{2}=\begin{cases}\alpha_{1}\mathrm{MSE}(\mathcal{M}_{d},I_{\mathrm{density}})+\alpha_{2}\mathcal{L}_{1},&y=0\\ \mathcal{L}_{1},&y=1\end{cases}$ (7)
where $\mathrm{MSE}()$ is the mean absolute error, and $\mathcal{M}_{d}$ is obtained from $I_{\mathrm{haze}}$ and the learnable paired prompts $T=[T_{\mathrm{haze}},T_{\mathrm{clear}}]$ according to Equation 5. $\alpha_{1}$ and $\alpha_{2}$ are the weights of different loss functions. More details is in supplementary material. Training is lightweight since there is no need to fine-tune the CLIP image encoder.
Finally, we use the learned paired prompts $[T_{\mathrm{haze}},T_{\mathrm{clear}}]$ to generate the estimated patch-wise haze density map $\mathcal{M}_{d}$. Some examples are shown in Figure 5, demonstrating that our method is an effective and universal approach applicable to both homogeneous and non-homogeneous haze.
3.3.2 Obtaining Refined Semantic Map
Considering that the pretrained image encoder of CLIP is obtained through a classification pretext task, each location in the feature map before pooling captures regional semantic information [74]. Therefore, we reuse $F_{\mathrm{image}}$ from Equation 4 and regard it as semantic information. This approach not only provides high-level information but also avoids introducing additional computational overhead.
However, relying solely on $F_{\mathrm{image}}$ leads to suboptimal performance, primarily because it is an abstract representation and lacks low-level semantic details. To address this limitation, we incorporate the input features of the TrambaDA blocks, $F_{\mathrm{input}}$, to supplement $F_{\mathrm{image}}$ and introduce a bidirectional cross-attention mechanism to align their semantic information across different scales, thereby yielding a refined semantic map $\mathcal{M}_{s}$:
$\begin{split}\mathcal{M}_{s}=\mathrm{Linear}&(\text{Attention}_{\text{input}\to\text{image}}(Q_{image},K_{input},V_{input}),\\ &(\text{Attention}_{\text{image}\to\text{input}}(Q_{input},K_{image},V_{image})).\end{split}$ (8)
$Q_{i}$, $K_{i}$ and $V_{i}$ represent the query, key, and value derived from $F_{\mathrm{i}}$ after channel reduction or adaptive pooling, where $\mathrm{i}\in\{\mathrm{image},\mathrm{input}\}$. This process is shown in Figure 4.
3.3.3 Generating Dual-path Aggregation Weights
Finally, the estimated haze density map $\mathcal{M}_{d}$ and the obtained refined semantic map $\mathcal{M}_{s}$ are jointly used to generate the dual-path aggregation weights, as follows:
$\begin{split}W_{short},W_{long}&=\mathrm{Split}(\mathrm{Linear}(\mathrm{Interpolate}(\\ &\quad\quad\quad\quad\mathrm{Concat}(\mathcal{M}_{d},\mathcal{M}_{s})))).\end{split}$ (9)
Here, $\mathrm{Concat}(\cdot)$ denotes a concatenation operation, $\mathrm{Interpolate}(\cdot)$ represents an interpolation operation used to resize $\mathcal{M}_{d}$ and $\mathcal{M}_{s}$ to match the dimensions of $F_{\mathrm{out\_{short}}}$ and $F_{\mathrm{out\_{long}}}$, $\mathrm{Linear}(\cdot)$ refers to a projection layer, and $\mathrm{Split}(\cdot)$ indicates a channel-wise splitting operation. The aggregation is subsequently performed using Equation 3.
4 Experiments
Figure 6: Visual comparisons on the NH-Haze dataset [3]. The red boxes highlight the restoration effects for thick haze and texture details. Please zoom in for the best view.
Figure 7: Visual comparisons on RESIDE SOTS Indoor dataset [38]. The red boxes and green boxes show the restoration of high-frequency and low-frequency information, respectively.
4.1 Experiment Setup
Training Details. DehazeMatic is implemented with PyTorch on NVIDIA A100 GPUs. We use Adam [32] optimizer with default parameters ($\beta_{1}=0.9$, $\beta_{2}=0.99$) and a cosine annealing strategy [46] with restarts. The initial learning rate is set to $2\times 10^{-4}$, gradually decreasing to $2\times 10^{-6}$. We train the homogeneous haze dataset for 200 epochs and the non-homogeneous haze dataset for 400 epochs. The images are randomly cropped to a size of 256$\times$256 and augmented with flipping. We use L1 and perceptual loss [29] to supervise dehazing process.
Datasets. We evaluate the proposed DehazeMatic on homogeneous (i.e., RESIDE [38]) and non-homogeneous haze datasets (i.e., NH-HAZE [3] and Dense-Haze [2]). The RESIDE [38] includes two training subsets: the Indoor Training Set (ITS) with 13,990 pairs of indoor images, and the Outdoor Training Set (OTS) with 313,950 pairs of outdoor images. Models are evaluated on the corresponding subset of the Synthetic Objective Testing Set (SOTS). The NH-HAZE [3] and Dense-Haze [2] use a professional haze generator to simulate real hazy conditions, making them suitable for evaluating model performance in real-world non-homogeneous haze removal. These datasets contain 55 and 55 paired images, respectively. We use the last 5 images as the testing set and the remaining images as the training set.
4.2 Comparison with SOTA Methods
Table 1 shows the quantitative results on various benchmarks. As the parameters of the CLIP encoder are frozen, we exclude them from the overhead, which is consistent with previous work. Notably, the additional inference time introduced by CLIP is minimal, requiring only 3.8 ms per image on an NVIDIA A100 GPU.
Performance on Non-homogeneous Haze. MITNet [58] and OKNet [10] are competitive in terms of PSNR and SSIM, respectively; however, our DehazeMatic achieves the highest performance in both metrics across all datasets. Moreover, as shown in Figure 6, MITNet [58] and OKNet [10] suffer from blue artifacts and blurred details, respectively. In contrast, DehazeMatic generates haze-free images with optimal fidelity and richer structural details while maintaining the highest quantitative scores.
Figure 8: Illustration of trade-off between performance and overhead. The line chart in (a) represents the average PSNR on the SOTS-Indoor and SOTS-Outdoor datasets, while the line chart in (b) represents the average PSNR on the NH-Haze and Dense-Haze datasets. The greater the extent to which the line chart’s value exceeds that of the bar chart, the better the model’s trade-off.
Table 2: Quantitative comparisons of various methods on real-world haze dataset RTTS [38]. The best results and the second best results are in bold and underline, respectively.
No-reference Metrics FFA-Net MB-TaylorFormer-B MITNet DehazeFormer DEA-Net OKNet DCMPNet DehazaMatic
[…]
ENTROPY $\uparrow$ 7.164 6.907 7.080 7.090 7.079 7.112 7.143 7.405
Performance on Homogeneous Haze. DehazeMatic achieves the SOTA in the average performance of indoor and outdoor scenes, and especially significantly outperforms the second-best method by 0.79 dB with fewer MACs in outdoor scenes. Figure 7 provides visual comparisons on the SOTS Indoor dataset [38]. In the green and red boxed regions, DehazeMatic exhibits superior color consistency and effectively restores finer details, respectively.
Trade-off between Performance and Overhead. As shown in Figure 8, our DehazeMatic achieves the best trade-off for both types of haze, demonstrating its high efficiency.
Generalization on Real-world Haze Dataset. To evaluate generalization, we compare DehazeMatic with several competitive methods on the real-world haze dataset RTTS [38], using models trained on the RESIDE OTS dataset [38] for inference. As shown in Table 2 , our method outperforms all others across no-reference metrics, highlighting its superior generalization ability and practical applicability.
4.3 Ablation Studies
We conduct extensive ablation studies to demonstrate the effectiveness of each component. For a fair comparison, we tune the hyperparameters of variant models to ensure that their overhead matches that of DehazeMatic.
Table 3: Ablation study of the explicit coexistence of short- and long-range dependencies.
Setting SOTS-Indoor NH-Haze
$\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$
Only Short Dependencies 38.41 0.992 20.40 0.771
Only Long Dependencies 39.78 0.994 20.56 0.779
DehazeMatic 41.50 0.996 21.47 0.806
Only Convolution 38.29 0.992 20.33 0.767
Only Linear Attention 40.06 0.994 20.65 0.780
Conv. & Linear Attention 41.41 0.996 21.36 0.803
Table 4: Ablation study of CLIP-enhanced Dual-path Aggregator.
Setting SOTS-Indoor NH-Haze
$\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$
(a) Remove CedA Addition 39.55 0.993 20.71 0.781
Concatenation 39.80 0.994 20.74 0.784
(b) Overall Design W/o Density Map 40.88 0.995 20.91 0.792
W/o Semantic Map 41.02 0.995 21.23 0.794
Transmission Map 40.97 0.995 21.27 0.797
(c) Density Map Depth Map 41.16 0.996 21.28 0.795
Predefined Prompts 40.24 0.994 20.82 0.789
(d) Sematic Map W/o High-level 41.12 0.995 21.30 0.800
W/o Low-level 41.19 0.996 21.34 0.802
DehazeMatic 41.50 0.996 21.47 0.806
Explicit Coexistence of Short- and Long-range Dependencies. We remove either the Mamba path (i.e., long-range) or the Transformer path (i.e., short-range) in the TrambaDA blocks to construct variant models, while simultaneously excluding the CedA module. The corresponding results are shown in the first three rows of Table 3.
Additionally, to verify that its superiority is not dependent on a specific feature extractor, we use convolution [78] (i.e., short-range), linear self-attention [51] (i.e., long-range) and the proposed CedA module to construct a variant model and repeat the above experiments. The corresponding quantitative results are presented in the last three rows of Table 3, demonstrating the versatility of our idea. Although linear self-attention [51] achieves better results than the adopted Mamba block, its time cost is significantly larger, so we do not adopt it. The comprehensive results in Table 3 demonstrate that the complementary short- and long-range dependencies introduced by the dual-path design, along with their reasonable aggregation, can significantly enhance image dehazing.
Figure 9: Visual comparisons of haze density maps estimated using learned prompts and manually predefined prompts.
CLIP-enhanced Dual-path Aggregator (CedA). As shown in Table 4, we conduct extensive ablation studies on the CedA to verify the effectiveness of each component.
(a) We replace the CedA with addition and concatenation operations to verify the necessity of adaptively aggregating short- and long-range dependencies in the dual paths.
(b) We then verify the importance of jointly using haze density and semantic information for aggregation guidance by removing each component individually.
(c) Next, we investigate the estimated haze density map. First, we verify the rationale for using the density map rather than other commonly used guidance signals in dehazing (i.e., the transmission map from DCP [22] and the depth map from Depth Anything [67]) to guide the aggregation. The results indicate that haze density is a more suitable influencing factor.
Furthermore, we use manually predefined prompts to estimate the haze density map to verify the effectiveness of using learnable prompts. As shown in Table 4, predefined prompts lead to a significant performance drop, even falling behind the approach that relies solely on semantic maps for aggregation guidance. In addition, as illustrated in Figure 9, predefined prompts are entirely incapable of generating valid patch-wise haze density map.
(d) Finally, we validate the necessity of each type of information in the refined semantic map by ablating either high-level semantic information (i.e., from CLIP) or low-level semantic information (i.e., input of Tramba blocks).
5 Conclusion
In this paper, we propose DehazeMatic, a framework that employs a dual-path design to explicitly and simultaneously capture complementary local fine-grained details and global contextual cues. Based on key insights from our experiments—namely that different tokens in an image require varying ranges of dependencies and that haze density and semantic information are the primary factors influencing the required range—we introduce the CLIP-enhanced Dual-path Aggregator. It leverages the rich semantic priors and powerful generalization ability encapsulated in CLIP to ensure that each dependency contributes optimally to performance while further exploring CLIP’s potential in image dehazing. Extensive experiments demonstrate that DehazeMatic outperforms SOTA methods on various benchmarks and holds promise for inspiring future research.

### v3

@@
 Haze removal aims to restore a clear image from a hazy input.
-Existing methods have shown significant efficacy by capturing either short-range dependencies for local detail preservation or long-range dependencies for global context modeling.
-Given the complementary strengths of both approaches, a intuitive advancement is to explicitly integrate them into a unified framework.
-However, this potential remains underexplored in current research.
-In this paper, we propose DehazeMatic, which leverages the proposed Transformer-Mamba Dual Aggregation block to simultaneously and explicitly captures both short- and long-range dependencies through dual-path design for improved restoration.
-To ensure that dependencies at varying ranges contribute optimally to performance, we conduct extensive experiments to identify key influencing factors and determine that an effective aggregation mechanism should be guided by the joint consideration of haze density and semantic information.
-Building on these insights, we introduce the CLIP-enhanced Dual-path Aggregator, which utilizes the rich semantic priors encapsulated in CLIP and the estimated haze density map, derived from its powerful generalization ability, to instruct the aggregation process.
-Extensive experiments demonstrate that DehazeMatic outperforms sort-of-the-art methods across various benchmarks.
-Figure 1: Illustration of factors influencing the relative importance between short-range and long-range dependencies.
-We train a Transformer-based dehazing network with a global receptive field and synthesis hazy images with different densities based on the Atmospheric Scattering Model [49].
-(a) The horizontal axis represents the Euclidean distance between the current token and all other tokens, while the vertical axis represents the average attention weights between the current token and all tokens at the corresponding distance.
-The curve data is derived from all tokens of all images in the test set.
-(b) The bottom visualizes the attention map between the chosen tokens (indicated by the tails of red arrows) and all other tokens.
-On the left is the visualization of tokens at the same position corresponding to the same ground truth image, but with different haze densities.
-On the right is a visualization of tokens with different semantic information on the same homogeneous haze image.
-More analysis is in supplementary material.
-1 Introduction
-Single image dehazing aims to recover a clear image from a hazy input, serving as a crucial pre-processing step for high-level vision tasks under hazy conditions, such as object detection [39] and semantic segmentation [53].
-Figure 2: Illustration of CLIP [52]’s potential to perceive haze and its density.
-We use the ViT-B/32 OpenCLIP [26] model, pre-trained on the LAION dataset, for our experiments.
+Existing methods achieve notable success by specializing in either short-range dependencies to preserve local details or long-range dependencies to capture global context.
+Given the complementary strengths of both, a natural progression is to explicitly integrate them within a unified framework and enable their reasonable aggregation.
+However, this integration remains underexplored.
+In this paper, we propose DehazeMatic, which simultaneously and explicitly captures both short- and long-range dependencies through a dual-stream design.
+To optimize the contribution of dependencies at varying ranges, we conduct extensive experiments to identify key influencing factors and find that an effective aggregation mechanism should be guided by the joint consideration of haze density and semantic information.
+Building on these insights, we introduce the CLIP-enhanced Dual-path Aggregator, which not only enables the generation of fine-grained haze density maps for the first time, but also produces semantic maps within a shared backbone, ultimately leveraging both to instruct the aggregation process.
+Extensive experiments demonstrate that DehazeMatic outperforms state-of-the-art methods across multiple benchmarks.
+Introduction
+Image dehazing aims to recover a clear image from a hazy input, serving as a crucial pre-processing step for high-level vision tasks in hazy conditions, such as object detection (Li et al.
+2023) and semantic segmentation (Ren et al.
+2024).
+Current data-driven methods can be categorized into two classes based on the receptive field size of the operators used for feature extraction, namely (i) using convolution (Bai et al.
+2022; Cai et al.
+2016; Dong et al.
+2020; Li et al.
+2017; Ren et al.
+2018, 2020; Zhang and Patel 2018) or window-based self-attention (Kulkarni, Phutke, and Murala 2022; Kulkarni and Murala 2023; Song et al.
+2023; Wang et al.
+2023) and (ii) using linear self-attention (Qiu et al.
+2023) or State Space Model (SSM) (Shen et al.
+2023; Zhou et al.
+2024; Zhang et al.
+2024a).
+The former captures short-range dependencies, which provide fine-grained local perception but lack effective global modeling capabilities (Kim et al.
+2023; Veit, Wilber, and Belongie 2016; De and Smith 2020).
+In contrast, the latter captures long-range dependencies with rich contextual information while ensuring efficient computation, yet fails to preserve local 2D inductive biases (Huang et al.
+2024; Dosovitskiy 2020).
+Both types of methods have achieved excellent performance, and their respective strengths can compensate for each other’s limitations.
+This naturally motivates the design of a dual-stream network that explicitly integrates short- and long-range dependencies, thereby harnessing the complementary advantages of both approaches and potentially achieving superior performance.
+However, reasonably aggregating short- and long-range dependencies from these two paths remains a non-trivial task, as tokens with different characteristics in the image require varying ranges of dependencies.
+Simple operations such as addition, concatenation, or gating mechanisms (Ren et al.
+2018; Zhang et al.
+2020a; Hou et al.
+2025) often lead to suboptimal performance.
+More importantly, there is still no clear guidance on how to determine the relative importance between short-range and long-range dependencies for each token.
+In this paper, we are the first to identify, through both quantitative and qualitative experiments, the key factors that determine the relative importance of short- and long-range dependencies required by each token: haze density and semantic information.
+Based on these insights, we propose a dual-stream framework for dehazing, termed DehazeMatic, which simultaneously captures both short- and long-range dependencies.
+Since feature extraction operators are not the primary focus of this work, their selection is relatively flexible.
+We adopt window-based self-attention (Liu et al.
+2021) and the Selective Scan Space State Sequential Model (S6) block (Gu and Dao 2023) as representative implementations.
+Figure 1: Illustration of CLIP (Radford et al.
+2021)’s potential to perceive haze and its density.
+We use the ViT-B/32 OpenCLIP (Ilharco et al.
+2021) model, pre-trained on the LAION dataset, for our experiments.
 The values above the images indicate the PSNR and SSIM metrics, which reflect haze density from a quantitative perspective.
@@
 Notably, as haze density increases, the similarity score with the haze-describing prompt also increases.
-Current data-driven methods can be categorized into two classes based on the receptive field size of the operators used for feature extraction, namely (i) using convolution [5, 7, 13, 37, 55, 56, 72] or window-based self-attention [35, 34, 59, 64] and (ii) using linear self-attention [51] or State Space Model (SSM) [58, 80, 73].
-The former excels at capturing short-range dependencies but lacks effective global modeling capabilities [30, 62, 12].
-In contrast, the latter directly models long-range dependencies while ensuring efficient computation; however, it fails to preserve local 2D inductive bias [25, 14].
-Both categories of methods have achieved excellent performance, and their respective strengths can address the limitations of the other.
-Therefore, a natural idea is that explicitly combining short- and long-range dependencies (i.e., leveraging the strengths of both approaches) is expected to yield even more promising results.
-However, this integration has not been well explored.
-To address this gap, we will employ a dual-path design that concurrently captures fine-grained perception from nearest neighbor features and holistic contextual cues from non-local relationships.
-The ensuing problem is that reasonably integrating short- and long-range dependencies from these two paths is a non-trivial task, because tokens with different characteristics in the image require varying ranges of dependencies.
-Simple addition or concatenation often leads to suboptimal performance.
-More importantly, there is still no clear guidance on how to determine the relative importance between short-range and long-range dependencies for each token.
-One preliminary hypothesis is that haze density serves as a determining factor.
-To verify this, we conduct an experiment.
-As shown in Figure 1(a), higher haze density increases reliance on long-range dependencies, whereas lower haze density requires more short-range dependencies.
-Furthermore, additional investigation reveals that tokens with varying semantic information require different ranges of dependencies, as illustrated on the right side of Figure 1(b).
-Based on these findings, we conclude that haze density and semantic information jointly determine the relative importance of short-range and long-range dependencies.
-In this paper, building on the above insights, we propose DehazeMatic, where Mamba and Transformer are Integrated by CLIP for image dehazing.
-The core building block of the DehazeMatic is the proposed Transformer-Mamba Dual Aggregation (TrambaDA) block, which employs a dual-path design.
-In one path, we utilize window-based self-attention [45] to capture short-range dependencies.
-In the other path, inspired by recent advances in SSMs [18, 19, 44, 82, 42], we incorporate the Selective Scan Space State Sequential Model (S6) block [18], which maintains global information indefinitely with linear complexity.
-We also adopt the four-way scanning strategy to mitigate the “direction-sensitive" problem caused by the non-causal nature of visual data [44].
-Next, to ensure that both short- and long-range dependencies contribute optimally to performance, a dual-path aggregator capable of adaptively learning the relative importance of each path is required.
-As indicated by Figure 1, the aggregator must capture both the haze density and semantic information of each token in the input image.
-Recent advances in CLIP [52], pre-trained on web-scale datasets, reveal that it not only encapsulates rich semantic priors (e.g., enabling strong performance in zero-shot semantic segmentation [81, 76]), but also has the potential to perceive haze and its density (as shown in Figure 2).
-Motivated by this, we propose the CLIP-enhanced Dual-path Aggregator (CedA).
-Specifically, we treat the affinity scores between the the patch-wise image embeddings from the CLIP image encoder and the text embeddings of paired prompts describing clear and hazy conditions from the CLIP text encoder as the estimated haze density map.
-To find the best prompts for better estimation and avoid labor-intensive prompt engineering, these paired prompts are learned.
-For semantic information, we reuse the the patch-wise image embeddings to provide high-level and abstract features while avoiding additional computational overhead.
-To supplement the low-level details, we incorporate the input features from the TrambaDA blocks and align both representations using a bidirectional cross-attention mechanism.
-Finally, we generate dual-path aggregation weights based on the refined semantic information and the estimated haze density map.
-Figure 3: (a) The architecture of our proposed DehazeMatic.
-(b) The inner structure of Transformer-Mamba Dual Aggregation (TrambaDA) block.
-The weights generated by CLIP-enhanced Dual-path Aggregator (CedA) are used to aggregate the outputs of dual-path.
-Our contributions can be summarized as follows:
-  • We propose DehazeMatic, which demonstrates the superiority of explicitly and simultaneously capturing both short- and long-range dependencies while effectively aggregating dependencies across varying ranges.
-It achieves SOTA performance in various benchmarks.
-  • For the first time, we identify the key factors required for reasonably aggregating dependencies across different ranges in image dehazing.
-Based on these insights, we propose the CedA module, which performs dynamic aggregation according to haze density and semantic characteristics of each token within the image.
-  • The proposed CedA module further explore the potential of CLIP in image dehazing and, for the first time, demonstrate its capability in haze density estimation, all without the need for fine-tuning the encoder.
-2 Related Work
-2.1 Single Image Dehazing
+Next, based on the identified key factors, we aim to generate a haze density map and a semantic information map to guide the aggregation of dependencies across varying ranges.
+Moreover, to improve model efficiency, we seek to generate both maps using a shared backbone.
+However, this objective presents significant challenges: not only is there currently no method capable of accurately estimating haze density maps, but extracting two types of information with substantial modality differences from a single backbone remains inherently difficult.
+Inspired by recent advances in CLIP (Radford et al.
+2021), which is pretrained on web-scale datasets and encapsulates rich semantic priors that enable strong performance in zero-shot semantic segmentation (Zhou et al.
+2023; Zhang et al.
+2024b), we further observe that CLIP has the potential to perceive haze and its density, as demonstrated in Figure 1.
+Building on this observation, we propose CLIP-enhanced Dual-path Aggregator (CedA).
+Specifically, we treat the patch-wise image embeddings from the CLIP image encoder as the semantic map.
+In parallel, we compute affinity scores between these image embeddings and the text embeddings of paired prompts describing clear and hazy conditions—obtained from the CLIP text encoder—and use these scores as the estimated haze density map.
+To enable accurate estimation while avoiding labor-intensive prompt engineering, we replace manually designed prompts with learnable ones.
+Training proceeds in a progressive two-stage manner: we first employ cross-entropy loss, followed by regression-based training on a self-constructed triplet dataset.
+Finally, based on the generated haze density and semantic maps, the dual-path aggregation weights are computed to reasonably and effectively aggregate short- and long-range dependencies.
+Our contributions are summarized as follows:
+  • We propose DehazeMatic, showing that once effective aggregation is achieved, explicitly and simultaneously modeling short- and long-range dependencies leads to state-of-the-art performance on multiple benchmarks.
+  • We are the first to identify the key factors that determine the relative importance of short- and long-range dependencies in image dehazing.
+Building on this insight, we propose the CedA module, which dynamically aggregates these dependencies without being restricted to any specific feature extraction operator.
+  • We further explore the potential of CLIP in haze removal and, for the first time, achieve accurate estimation of haze density maps, without any fine-tuning of the encoder.
+Related Work
+Single Image Dehazing
 Due to spatially variant transmission map and atmospheric light, single image dehazing is a highly ill-posed problem.
-Early prior-based methods [23, 16, 31, 60, 83, 6] achieved dehazing by imposing various assumptions to estimate key parameters in the Atmospheric Scattering Model (ASM) [49].
+Early prior-based methods (He et al.
+2016; Fattal 2008; Kim, Park, and Eom 2019; Tan 2008; Zhu, Mai, and Shao 2015; Berman, Treibitz, and Avidan 2018) achieved dehazing by imposing various assumptions to estimate key parameters in the Atmospheric Scattering Model (ASM) (Narasimhan and Nayar 2002).
 ASM provides strong theoretical support and interpretability for these methods; however, they often fail when images deviate from statistical laws.
-With the rapid advancement of deep learning [33], numerous learning-based methods have been proposed, resulting in improved performance.
-Early methods [7, 37] employ neural networks to estimate key parameters in the ASM and subsequently restore the haze-free images.
-Later, ASM-independent deep networks [54, 55, 43, 40, 57, 13, 75, 50, 36, 66, 70, 59, 17, 68, 77, 8, 15, 9, 65, 69, 11] directly estimate clear images or haze residuals.
-DehazeFormer [59] is a representative method that achieves efficient feature extraction through window-based self-attention and several key modifications.
+With the rapid advancement of deep learning (Krizhevsky, Sutskever, and Hinton 2012), various learning-based methods have been proposed, resulting in improved performance.
+Early methods (Cai et al.
+2016; Li et al.
+2017) employ neural networks to estimate key parameters in the ASM and subsequently restore the haze-free images.
+Later, ASM-independent deep networks (Ren et al.
+2016, 2018; Liu et al.
+2019; Li et al.
+2019; Shao et al.
+2020; Dong et al.
+2020; Zhang et al.
+2020b; Qin et al.
+2020; Li et al.
+2020; Wu et al.
+2021; Ye et al.
+2022; Song et al.
+2023; Feng et al.
+2024; Yang et al.
+2024b; Zhang, Zhou, and Li 2024; Chen, He, and Lu 2024; Fang et al.
+2024b; Cong et al.
+2024; Wang et al.
+2024b; Yang et al.
+2025; Cui et al.
+2025) directly estimate clear images or haze residuals.
+DehazeFormer (Song et al.
+2023) is a representative method that achieves efficient feature extraction through window-based self-attention and several key modifications.
 However, its inherently limited receptive field constrains its performance potential.
-To achieve a global receptive field with lower computational overhead, some methods [58, 71] incorporate frequency domain features into image dehazing, offering the additional benefit of more direct learning of high-frequency information.
-For the same purpose, other researchers [79, 73] introduce the Mamba [18] architecture into image dehazing.
+To achieve a global receptive field with lower computational overhead, some methods (Shen et al.
+2023; Yu et al.
+2022) incorporate frequency domain features into image dehazing, offering the additional benefit of more direct learning of high-frequency information.
+For the same purpose, other researchers (Zheng and Wu 2024; Zhang et al.
+2024a) introduce the Mamba (Gu and Dao 2023) architecture into dehazing.
 However, none of these methods address how to better integrate global semantic features with local texture features.
-2.2 CLIP for Low-Level Vision Tasks
-Classic vision-language models, such as CLIP [52], aim to learn aligned features in the embedding space from image-text pairs using contrastive learning.
+CLIP for Low-Level Vision Tasks
+Classic vision-language models like CLIP (Radford et al.
+2021), aim to learn aligned features in the embedding space from image-text pairs using contrastive learning.
 Some studies have explored leveraging the rich prior knowledge encapsulated in CLIP to assist with low-level vision tasks.
-In All-in-One image restoration, some researchers [47, 1, 27] use degradation embeddings from the CLIP image encoder to implicitly guide the network in making adaptive responses.
-In monocular depth estimation, recent studies [74, 4, 24] use CLIP to map patches of the input image to specific semantic distance tokens, which are then projected onto a quantified depth bin for estimation.
-In low-light enhancement, some methods [41, 48] use text-image similarity between the enhanced results and learnable prompt pairs to train the enhancement network.
-Figure 4: Illustration of the CLIP-enhanced Dual-path Aggregator generating aggregation weights.
-To ensure computational efficiency, the TrambaDA blocks share the same aggregation weights at each stage.
-Some works also introduce CLIP into image [63] and video [53] dehazing, but they all use text-image similarity between dehazed results and contrastive prompt sets as regularization to guide the restoration process.
+In All-in-One image restoration, some researchers (Luo et al.
+2023; Ai et al.
+2024; Jiang et al.
+2025) use degradation embeddings from the CLIP image encoder to implicitly guide the network in making adaptive responses.
+In monocular depth estimation, recent studies (Zhang et al.
+2022; Auty and Mikolajczyk 2023; Hu et al.
+2024) use CLIP to map patches of the input image to specific semantic distance tokens, which are then projected onto a quantified depth bin for estimation.
+In low-light enhancement, some methods (Liang et al.
+2023; Morawski et al.
+2024) use text-image similarity between the enhanced results and learnable prompt pairs to train the enhancement network.
+Some works also introduce CLIP into image (Wang et al.
+2024a) and video (Ren et al.
+2024) dehazing, but they all use text-image similarity between dehazed results and contrastive prompt sets as regularization to guide the restoration process.
 In contrast, our method further exploits the potential of CLIP by directly incorporating its latent embeddings into the main network to guide the dehazing process.
-3 Method
-3.1 Overall Architecture
-The architecture of DehazeMatic is depicted in Figure 3.
-Given a hazy image, shallow features are first extracted and then passed through a symmetric encoder-decoder framework.
-Each stage of encoder and decoder consists of multiple Transformer-Mamba Dual Aggregation (TrambaDA) blocks, along with corresponding downsampling or upsampling operations.
-We also incorporate skip connections between the encoder and decoder at each level.
-Finally, the output of the last decoder stage is combined with the input hazy image to produce the haze-free image.
-3.2 Transformer-Mamba Dual Aggregation Block
-The TrambaDA block’s backbone employs a dual-path design, with one path capturing short-range dependencies and the other capturing long-range dependencies, thereby providing complementary restoration information for image tokens with distinct characteristics.
-Capturing Short-range Dependencies.
-We construct this path using window-based self-attention [45], which offers stronger fitting capability through dynamic weights.
-Specifically, given the input feature $F_{\rm{in}}\in\mathbb{R}^{H\times W\times C}$, we first divide $F_{\rm{in}}$ into non-overlapping windows with size of $M\times M$, then obtain flattened feature $F_{\mathrm{in}}^{i}\in\mathbb{R}^{M^{2}\times C}$ of each window $i$.
-For simplicity, we only consider the single-head attention.
-Performing self-attention on non-overlapping windows can be formulated as:
-$\begin{split}F_{\mathrm{out\_{short}}}^{i}=&\text{Attention}(F_{\mathrm{in}}^{i}W^{Q},F_{\mathrm{in}}^{i}W^{K},F_{\mathrm{in}}^{i}W^{V}),\\ \text{Attention}&(Q,K,V)=\text{Softmax}(\frac{{QK}^{T}}{\sqrt{d_{k}}})V,\\ F_{\mathrm{out\_{short}}}=&\{F_{\mathrm{out\_{short}}}^{1},F_{\mathrm{out\_{short}}}^{2},\cdots,F_{\mathrm{out\_{short}}}^{N}\},\end{split}$ (1)
-where $N=HW/M^{2}$; $W^{Q},W^{K},W^{V}\in\mathbb{R}^{C\times d_{k}}$ are projection matrices for queries, keys, and values, respectively.
-Figure 5: Illustration of estimated haze density maps.
-The proposed method can identify the relative haze density not only in different regions of non-homogeneous haze images but also between two homogeneous haze images.
-Capturing Long-range Dependencies.  We build this path using Mamba [18]’s S6 block, which preserves the global receptive field with linear complexity and incorporates hardware-aware algorithms.
-Due to the “direction-sensitive” problem caused by the non-causal nature of visual data, directly applying the S6 block on the flattened image will result in suboptimal performance [44].
-Following Vmamba [44], we unfold feature map along four directions, generating four independent sequences.
-Each feature sequence is processed through the S6 block, and all output sequences are merged into a new sequence.
-This process can be represented as follows:
-$\begin{split}F_{\mathrm{in}}^{d}=expand&(F_{\mathrm{in}},d),d=\{1,2,3,4\},\\ \bar{F}&{}_{\mathrm{in}}^{d}=\mathrm{S6}(F_{\mathrm{in}}^{d}),\\ F_{\mathrm{out\_{long}}}=m&erge(\bar{F}_{\mathrm{in}}^{1},\bar{F}_{\mathrm{in}}^{2},\bar{F}_{\mathrm{in}}^{3},\bar{F}_{\mathrm{in}}^{4}),\end{split}$ (2)
-where $expand(\cdot)$ and $merge(\cdot)$ correspond to the scan expand and scan merge operations, respectively.
-3.3 CLIP-enhanced Dual-path Aggregator
-The CLIP-enhanced Dual-path Aggregator (CedA) utilizes the generated pixel-level weights to adaptively aggregate dual-path outputs, ensuring that dependencies across varying ranges contribute optimally to performance:
+Motivational Experiment
+Optimal aggregation of dependencies is essential for a dual-stream network to fully leverage short- and long-range cues for dehazing.
+To this end, we empirically investigate the key factors that govern their relative importance, thereby enabling more reasonable and effective aggregation.
+Quantifying the Importance of Dependencies
+Dependency denotes the influence exerted by other tokens on the current token (Bengio, Simard, and Frasconi 1994; Hochreiter and Schmidhuber 1997), and its range can be quantified by the Euclidean distance.
+For the experiment presented in this section, we train a Transformer model with a global receptive field, where the attention weights assigned by other tokens to the current token in the self-attention mechanism are defined as the importance of the corresponding dependencies.
+Experimental Design
+We begin by hypothesizing potential key influencing factors and then verify their validity through quantitative and qualitative results.
+To this end, we select image tokens with differing characteristics in terms of the hypothesized factors as the current tokens, and examine whether the dependency importance at a fixed distance varies accordingly.
+This setup enables us to reflect the relative importance between short-range and long-range dependencies, as the importance of all other tokens with respect to the current token is normalized using L1 normalization.
+Experimental Observations
+Haze density is intuitively regarded as a key contributing factor.
+To validate this, we synthesize hazy images with varying density levels using the Atmospheric Scattering Model (Narasimhan and Nayar 2002) and conduct corresponding experiments.
+As illustrated in Figure 2, as the haze becomes denser, the relative importance of long-range dependencies increases, while that of short-range dependencies decreases—and vice versa.
+Quantitative results further confirm this observation.
+Semantic information is also hypothesized to be an influential factor, as prior work (Huang et al.
+2020) indicates that scenes with different levels of complexity require dependencies at varying ranges.
+To examine this, we conduct experiments on indoor and outdoor images from the RESIDE dataset (Li et al.
+2018), which generally exhibit distinct semantic characteristics.
+As shown in Figure 2, the quantitative results support our hypothesis, while qualitative results further demonstrate that for tokens with different semantic content within the same image, the relative importance of dependencies at different ranges also differs.
+Based on these findings, we are the first to propose that in image dehazing, the relative importance of short-range and long-range dependencies is jointly influenced by both haze density and semantic information.
+Figure 2: Illustration of factors influencing the relative importance between short-range and long-range dependencies.
+The left side presents how the quantitative results vary with changes in the characteristics of the hypothesized factors.
+The horizontal axis represents the Euclidean distance between the current token and all other tokens, i.e., the range of dependency, while the vertical axis indicates the average importance of other dependencies located at the corresponding distance to the current token.
+The curve is derived from all tokens across all images in the test set.
+The right side visualizes the importance of other tokens to the chosen tokens (marked with pentagrams).
+In (a), the tokens are located at the same position and correspond to the same ground truth image but exhibit different haze densities.
+In (b), the tokens differ in semantic content but are located within the same image under homogeneous haze conditions.
+Method
+Figure 3: (a) The overall architecture of DehazeMatic.
+(b) The internal structure of the Transformer-Mamba Dual Aggregation (TrambaDA) block.
+(c) The process by which the CLIP-enhanced Dual-path Aggregator (CedA) generates the aggregation weights used to fuse the dual-path outputs of the TrambaDA block.
+To improve computational efficiency, the TrambaDA blocks share the same aggregation weights within each stage.
+We propose DehazeMatic, as illustrated in Figure 3.
+It adopts a symmetric encoder-decoder framework in which each stage is constructed using a dual-stream design.
+To effectively aggregate the outputs from the two streams, we introduce the CLIP-enhanced Dual-path Aggregator, which utilizes the estimated haze density and semantic maps to instruct the aggregation process.
+Transformer-Mamba Dual Aggregation Block
+The TrambaDA block adopts a dual-stream design to simultaneously capture dependencies at different ranges, aiming to provide complementary restoration information for image tokens with distinct characteristics.
+The specific feature extractors for each path are flexible and not the focus of this paper.
+Therefore, for the sake of design simplicity, we adopt window-based self-attention  (Liu et al.
+2021) to capture short-range dependencies, with its output denoted as $F_{\mathrm{out\_{short}}}$.
+To model long-range dependencies, we apply Mamba’s S6 block  (Gu and Dao 2023), equipped with a four-directional scanning strategy (Liu et al.
+2024), and denote its output as $F_{\mathrm{out\_{long}}}$.
+The detailed structure is illustrated in Figure 3(b).
+CLIP-enhanced Dual-path Aggregator
+Based on the aforementioned insights, CedA is designed to first generate fine-grained haze density and semantic maps, and then produce pixel-level weights accordingly to adaptively aggregate the dual-path outputs, i.e.,
 $\begin{split}F_{out}=W_{short}F_{\mathrm{out\_{short}}}+W_{long}F_{\mathrm{out\_{long}}}.\end{split}$ (1)
-$W_{short},W_{long}\in\mathbb{R}^{H\times W}$ are the aggregation weights corresponding to short- and long-range dependencies, respectively, and their generation process is shown in Figure 4.
-3.3.1 Estimating Haze Density Map
-Define Estimation Pipeline.
+$W_{short},W_{long}\in\mathbb{R}^{H\times W}$ are the aggregation weights corresponding to short- and long-range dependencies, respectively, and their generation process is shown in Figure 3(c).
+Estimating Haze Density Map
+We begin by defining the estimation pipeline.
 Inspired by Figure 1, we leverage CLIP to estimate the haze density map.
@@
 where $\mathrm{sim}(\cdot,\cdot)$ is the similarity calculation, and final haze density map $\mathcal{M}_{d}\in\mathbb{R}^{H_{p}\times W_{p}}$.
-Table 1: Quantitative comparisons of various methods on various dehazing benchmarks. The best results and the second best results are in bold and underline, respectively.
-Methods Publication RESIDE SOTS (Homogeneous Haze) Non-homogeneous Haze Overhead
-Outdoor Indoor Average NH-Haze Dense-Haze
-$\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$Param MACs
-DCP [22] TPAMI-10 19.14 0.861 16.61 0.855 17.88 0.858 10.57 0.520 10.06 0.385 - -
-AOD-Net [37] ICCV-17 24.14 0.920 20.51 0.816 22.33 0.868 15.40 0.569 13.14 0.414 1.76K 0.12G
-GridDehazeNet [43] ICCV-19 30.86 0.982 32.16 0.984 31.51 0.983 18.33 0.667 14.96 0.533 0.96M 21.55G
-FFA-Net [50] AAAI-20 33.57 0.984 36.39 0.989 34.98 0.987 19.87 0.692 16.09 0.503 4.46M 288.86G
-DeHamer [20] CVPR-22 35.18 0.986 36.63 0.988 35.91 0.987 20.66 0.684 16.62 0.560 132.45M 59.25G
-SRDefog [28] ACCV-22 - - - - - - 20.99 0.610 16.67 0.500 12.56M 24.18M
-MAXIM-2S [61] CVPR-22 34.19 0.985 38.11 0.991 36.15 0.988 - - - - 14.10M 216.00G
-SGID-PFF [5] TIP-22 30.20 0.975 38.52 0.991 34.36 0.983 - - - - 13.87M 156.67G
-PMNet [70] ECCV-22 34.74 0.985 38.41 0.990 36.58 0.988 20.42 0.730 16.79 0.510 18.90M 81.13G
-MB-TaylorFormer-B [51] ICCV-23 37.42 0.989 40.71 0.992 39.07 0.991 - - 16.66 0.560 2.68M 38.50G
-MITNet [58] MM-23 35.18 0.988 40.23 0.992 37.71 0.990 21.26 0.712 16.97 0.606 2.73M 16.42G
-DehazeFormer [59] TIP-23 34.29 0.983 38.46 0.994 36.38 0.989 20.31 0.761 16.66 0.595 4.63M 48.64G
-SCANet [21] CVPR-23 - - - - - - 19.52 0.649 15.35 0.508 2.39M 258.63G
-DEANet [8] TIP-24 36.03 0.989 40.20 0.993 38.12 0.991 20.84 0.801 16.73 0.602 3.65M 32.23G
-UVM-Net [79] Arxiv-24 34.92 0.984 40.17 0.996 37.55 0.990 - - - - 19.25M 173.55G
-OKNet [10] AAAI-24 35.45 0.992 37.59 0.994 36.52 0.993 20.29 0.800 16.85 0.620 4.42M 39.54G
-DCMPNet [77] CVPR-24 36.56 0.993 42.18 0.996 39.37 0.995 - - - - 18.59M 80.42G
-DehazeMatic - 38.21 0.995 41.50 0.996 39.86 0.996 21.47 0.806 17.28 0.629 4.58M 35.50G
-Train Learnable Haze/Clear Prompts.
-To find the best prompts for improved estimation and avoid time-consuming prompt engineering, we use learnable prompt tokens instead of manually predefined prompts to describe abstract haze and clear conditions.
-The training process consists of two stages.
-In the first stage, we optimize the learnable paired prompts using cross-entropy loss, enabling them to preliminarily distinguish between haze and clear conditions.
-Given a hazy and clear image $I_{\mathrm{haze}},I_{\mathrm{clear}}\in\mathbb{R}^{H\times W\times 3}$, we define the first-stage loss $\mathcal{L}_{1}$ as:
+To optimize estimation performance and reduce reliance on time-consuming prompt engineering, we employ learnable prompt tokens rather than manually predefined prompts to characterize abstract haze and clear conditions.
+The training comprises two stages.
+In Stage 1, we optimize learnable paired prompts using cross-entropy loss, enabling them to preliminarily distinguish between hazy and clear images.
+Given a hazy and clear image $I_{\mathrm{haze}},I_{\mathrm{clear}}\in\mathbb{R}^{H\times W\times 3}$, the first-stage loss $\mathcal{L}_{1}$ is defined as:
 $\begin{split}\mathcal{L}_{1}=-(y*\log(\hat{y})+(1-y)*\log(1-\hat{y})),\\ \hat{y}=\frac{e^{cos(\mathrm{Pool}(\Phi_{\mathrm{image}}(I)),\Phi_{\mathrm{text}}(T_{\mathrm{clear}}))}}{\sum_{i\in\{\mathrm{haze},\mathrm{clear}\}}e^{cos(\mathrm{Pool}(\Phi_{\mathrm{image}}(I)),\Phi_{\mathrm{text}}(T_{i}))}},\end{split}$ (4)
@@
 Since $\Phi_{\mathrm{image}}$ is the CLIP image encoder without the final pooling layer, we apply a global pooling layer $\mathrm{Pool}()$ to obtain the latent image embedding with a shape of $\mathbb{R}^{1\times d_{c}}$.
-In the second stage, our goal is to achieve better prediction of haze density.
+In Stage 2, the objective is to predict haze density more accurately.
 The most straightforward and effective optimization approach is regression.
-However, there is no ground-truth density map corresponding to the hazy image in the existing dataset.
-To address this, we construct triplet data $\{I_{\mathrm{haze}},I_{\mathrm{clear}},I_{\mathrm{density}}\}$ based on Atmospheric Scattering Model [49].
-The second-stage loss $\mathcal{L}_{2}$ is defined as:
+However, existing datasets lack ground-truth density maps corresponding to hazy images.
+To address this, we construct triplet data $\{I_{\mathrm{haze}},I_{\mathrm{clear}},I_{\mathrm{density}}\}$ based on Atmospheric Scattering Model (Narasimhan and Nayar 2002).
+The second-stage loss $\mathcal{L}_{2}$ is then defined as:
 $\mathcal{L}_{2}=\begin{cases}\alpha_{1}\mathrm{MSE}(\mathcal{M}_{d},I_{\mathrm{density}})+\alpha_{2}\mathcal{L}_{1},&y=0\\ \mathcal{L}_{1},&y=1\end{cases}$ (5)
@@
 $\alpha_{1}$ and $\alpha_{2}$ are the weights of different loss functions.
-More details is in supplementary material.
 Training is lightweight since there is no need to fine-tune the CLIP image encoder.
-Finally, we use the learned paired prompts $[T_{\mathrm{haze}},T_{\mathrm{clear}}]$ to generate the estimated patch-wise haze density map $\mathcal{M}_{d}$.
+Finally, we use the learned paired prompts $T_{\mathrm{haze}},T_{\mathrm{clear}}]$ to generate the estimated patch-wise haze density map $\mathcal{M}_{d}$.
 Some examples are shown in Figure 4, demonstrating that our method is an effective and universal approach applicable to both homogeneous and non-homogeneous haze.
-3.3.2 Obtaining Refined Semantic Map
-Considering that the pretrained image encoder of CLIP is obtained through a classification pretext task, each location in the feature map before pooling captures regional semantic information [74].
+Figure 4: Illustration of estimated haze density maps.
+The proposed method can identify the relative haze density not only in different regions of non-homogeneous haze images but also between two homogeneous haze images.
+Obtaining Refined Semantic Map
+Considering that the pretrained image encoder of CLIP is obtained through a classification pretext task, each location in the feature map before pooling captures regional semantic information (Zhang et al.
+2022).
 Therefore, we reuse $F_{\mathrm{image}}$ from Equation 2 and regard it as semantic information.
 This approach not only provides high-level information but also avoids introducing additional computational overhead.
-However, relying solely on $F_{\mathrm{image}}$ leads to suboptimal performance, primarily because it is an abstract representation and lacks low-level semantic details.
-To address this limitation, we incorporate the input features of the TrambaDA blocks, $F_{\mathrm{input}}$, to supplement $F_{\mathrm{image}}$ and introduce a bidirectional cross-attention mechanism to align their semantic information across different scales, thereby yielding a refined semantic map $\mathcal{M}_{s}$:
-$\begin{split}\mathcal{M}_{s}=\mathrm{Linear}&(\text{Attention}_{\text{input}\to\text{image}}(Q_{image},K_{input},V_{input}),\\ &(\text{Attention}_{\text{image}\to\text{input}}(Q_{input},K_{image},V_{image})).\end{split}$ (8)
+However, using only $F_{\mathrm{image}}$ gives suboptimal results, primarily because it is an abstract representation and lacks low-level semantic details.
+To fix this, we incorporate the input features of the TrambaDA blocks, $F_{\mathrm{input}}$, to supplement $F_{\mathrm{image}}$ and introduce a bidirectional cross-attention mechanism to align their semantic information across different scales, thereby yielding a refined semantic map $\mathcal{M}_{s}$:
+$\begin{split}\mathcal{M}_{s}=\mathrm{Linear}&(\text{Attention}(Q_{image},K_{input},V_{input}),\\ &(\text{Attention}(Q_{input},K_{image},V_{image})).\end{split}$ (6)
 $Q_{i}$, $K_{i}$ and $V_{i}$ represent the query, key, and value derived from $F_{\mathrm{i}}$ after channel reduction or adaptive pooling, where $\mathrm{i}\in\{\mathrm{image},\mathrm{input}\}$.
-This process is shown in Figure 4.
-3.3.3 Generating Dual-path Aggregation Weights
+$\text{Attention}(\cdot)$ is self-attention operation.
+Generating Dual-path Aggregation Weights
 Finally, the estimated haze density map $\mathcal{M}_{d}$ and the obtained refined semantic map $\mathcal{M}_{s}$ are jointly used to generate the dual-path aggregation weights, as follows:
 $\begin{split}W_{short},W_{long}&=\mathrm{Split}(\mathrm{Linear}(\mathrm{Interpolate}(\\ &\quad\quad\quad\quad\mathrm{Concat}(\mathcal{M}_{d},\mathcal{M}_{s})))).\end{split}$ (7)
-Here, $\mathrm{Concat}(\cdot)$ denotes a concatenation operation, $\mathrm{Interpolate}(\cdot)$ represents an interpolation operation used to resize $\mathcal{M}_{d}$ and $\mathcal{M}_{s}$ to match the dimensions of $F_{\mathrm{out\_{short}}}$ and $F_{\mathrm{out\_{long}}}$, $\mathrm{Linear}(\cdot)$ refers to a projection layer, and $\mathrm{Split}(\cdot)$ indicates a channel-wise splitting operation.
+Here, $\mathrm{Concat}(\cdot)$ is concatenation operation, $\mathrm{Interpolate}(\cdot)$ represents an interpolation operation used to resize $\mathcal{M}_{d}$ and $\mathcal{M}_{s}$ to match the dimensions of $F_{\mathrm{out\_{short}}}$ and $F_{\mathrm{out\_{long}}}$, $\mathrm{Linear}(\cdot)$ refers to a projection layer, and $\mathrm{Split}(\cdot)$ indicates a channel-wise splitting operation.
 The aggregation is subsequently performed using Equation 1.
-4 Experiments
-Figure 6: Visual comparisons on the NH-Haze dataset [3]. The red boxes highlight the restoration effects for thick haze and texture details. Please zoom in for the best view.
-Figure 7: Visual comparisons on RESIDE SOTS Indoor dataset [38]. The red boxes and green boxes show the restoration of high-frequency and low-frequency information, respectively.
-4.1 Experiment Setup
+Experiments
+Experiment Setup
 Training Details.
 DehazeMatic is implemented with PyTorch on NVIDIA A100 GPUs.
-We use Adam [32] optimizer with default parameters ($\beta_{1}=0.9$, $\beta_{2}=0.99$) and a cosine annealing strategy [46] with restarts.
+We use Adam (Kingma and Ba 2014) optimizer with default parameters ($\beta_{1}=0.9$, $\beta_{2}=0.99$) and a cosine annealing strategy (Loshchilov and Hutter 2016) with restarts.
 The initial learning rate is set to $2\times 10^{-4}$, gradually decreasing to $2\times 10^{-6}$.
 We train the homogeneous haze dataset for 200 epochs and the non-homogeneous haze dataset for 400 epochs.
-The images are randomly cropped to a size of 256$\times$256 and augmented with flipping.
-We use L1 and perceptual loss [29] to supervise dehazing process.
+The images are randomly cropped to a size of $\times256$ and augmented with flipping.
+We use L1 loss and perceptual loss (Johnson, Alahi, and Fei-Fei 2016) to supervise dehazing process.
 Datasets.
-We evaluate the proposed DehazeMatic on homogeneous (i.e., RESIDE [38]) and non-homogeneous haze datasets (i.e., NH-HAZE [3] and Dense-Haze [2]).
-The RESIDE [38] includes two training subsets: the Indoor Training Set (ITS) with 13,990 pairs of indoor images, and the Outdoor Training Set (OTS) with 313,950 pairs of outdoor images.
-Models are evaluated on the corresponding subset of the Synthetic Objective Testing Set (SOTS).
-The NH-HAZE [3] and Dense-Haze [2] use a professional haze generator to simulate real hazy conditions, making them suitable for evaluating model performance in real-world non-homogeneous haze removal.
-These datasets contain 55 and 55 paired images, respectively.
-We use the last 5 images as the testing set and the remaining images as the training set.
-4.2 Comparison with SOTA Methods
-Table 1 shows the quantitative results on various benchmarks.
-As the parameters of the CLIP encoder are frozen, we exclude them from the overhead, which is consistent with previous work.
-Notably, the additional inference time introduced by CLIP is minimal, requiring only 3.8 ms per image on an NVIDIA A100 GPU.
-Performance on Non-homogeneous Haze.
-MITNet [58] and OKNet [10] are competitive in terms of PSNR and SSIM, respectively; however, our DehazeMatic achieves the highest performance in both metrics across all datasets.
-Moreover, as shown in Figure 6, MITNet [58] and OKNet [10] suffer from blue artifacts and blurred details, respectively.
-In contrast, DehazeMatic generates haze-free images with optimal fidelity and richer structural details while maintaining the highest quantitative scores.
-Figure 8: Illustration of trade-off between performance and overhead.
-The line chart in (a) represents the average PSNR on the SOTS-Indoor and SOTS-Outdoor datasets, while the line chart in (b) represents the average PSNR on the NH-Haze and Dense-Haze datasets.
-The greater the extent to which the line chart’s value exceeds that of the bar chart, the better the model’s trade-off.
-Table 2: Quantitative comparisons of various methods on real-world haze dataset RTTS [38]. The best results and the second best results are in bold and underline, respectively.
-No-reference Metrics FFA-Net MB-TaylorFormer-B MITNet DehazeFormer DEA-Net OKNet DCMPNet DehazaMatic
+We evaluate DehazeMatic on both synthetic and real-world datasets.
+The synthetic datasets encompass both homogeneous and non-homogeneous haze scenarios.
+For homogeneous haze, we utilize RESIDE dataset (Li et al.
+2018), which includes two training subsets: Indoor Training Set (ITS), containing 13,990 paired indoor images, and Outdoor Training Set (OTS), containing 313,950 paired outdoor images.
+Evaluation is conducted on the corresponding subsets of the Synthetic Objective Testing Set (SOTS).
+For non-homogeneous haze, we use the NH-HAZE (Ancuti, Ancuti, and Timofte 2020) and Dense-Haze (Ancuti et al.
+2019) datasets, both generated using a professional haze generator to simulate real-world hazy conditions.
+Each of these datasets comprises 55 paired images, of which the last 5 are used for testing and the remaining 50 for training.
+For real-world evaluation, we use the RTTS dataset (Li et al.
+2018), which contains 4,322 real-world hazy images.
+Comparison with SOTA Methods
+Table 1: Quantitative comparisons of various methods on multiple synthetic dehazing benchmarks. The best results and the second best results are in bold and underline, respectively.
+Methods Homogeneous Haze Non-homogeneous Haze Overhead
+SOTS-Outdoor SOTS-Indoor Average NH-Haze Dense-Haze Average
 […]
+DCP (He, Sun, and Tang 2010) 19.14 0.861 16.61 0.855 17.88 0.858 10.57 0.520 10.06 0.385 10.32 0.453 - -
+AOD-Net (Li et al. 2017) 24.14 0.920 20.51 0.816 22.33 0.868 15.40 0.569 13.14 0.414 14.27 0.492 1.76K 0.12G
+GridDehazeNet (Liu et al. 2019) 30.86 0.982 32.16 0.984 31.51 0.983 18.33 0.667 14.96 0.533 16.65 0.600 0.96M 21.55G
+FFA-Net (Qin et al. 2020) 33.57 0.984 36.39 0.989 34.98 0.987 19.87 0.692 16.09 0.503 17.98 0.598 4.46M 288.86G
+DeHamer (Guo et al. 2022) 35.18 0.986 36.63 0.988 35.91 0.987 20.66 0.684 16.62 0.560 18.64 0.622 132.45M 59.25G
+SRDefog (Jin et al. 2022) - - - - - - 20.99 0.610 16.67 0.500 18.83 0.555 12.56M 24.18M
+MAXIM-2S (Tu et al. 2022) 34.19 0.985 38.11 0.991 36.15 0.988 - - - - - - 14.10M 216.00G
+SGID-PFF (Bai et al. 2022) 30.20 0.975 38.52 0.991 34.36 0.983 - - - - - - 13.87M 156.67G
+PMNet (Ye et al. 2022) 34.74 0.985 38.41 0.990 36.58 0.988 20.42 0.730 16.79 0.510 18.61 0.620 18.90M 81.13G
+MB-TaylorFormer-B (Qiu et al. 2023) 37.42 0.989 40.71 0.992 39.07 0.991 - - 16.66 0.560 - - 2.68M 38.50G
+MITNet (Shen et al. 2023) 35.18 0.988 40.23 0.992 37.71 0.990 21.26 0.712 16.97 0.606 19.12 0.659 2.73M 16.42G
+DehazeFormer (Song et al. 2023) 34.29 0.983 38.46 0.994 36.38 0.989 20.31 0.761 16.66 0.595 18.49 0.595 4.63M 48.64G
+SCANet (Guo et al. 2023) - - - - - - 19.52 0.649 15.35 0.508 17.44 0.579 2.39M 258.63G
+DEANet (Chen, He, and Lu 2024) 36.03 0.989 40.20 0.993 38.12 0.991 20.84 0.801 16.73 0.602 18.79 0.702 3.65M 32.23G
+UVM-Net (Zheng and Wu 2024) 34.92 0.984 40.17 0.996 37.55 0.990 - - - - - - 19.25M 173.55G
+OKNet (Cui, Ren, and Knoll 2024) 35.45 0.992 37.59 0.994 36.52 0.993 20.29 0.800 16.85 0.620 18.57 0.710 4.42M 39.54G
+DCMPNet (Zhang, Zhou, and Li 2024) 36.56 0.993 42.18 0.996 39.37 0.995 - - - - - - 18.59M 80.42G
+DehazeMatic 38.21 0.995 41.50 0.996 39.86 0.996 21.47 0.806 17.28 0.629 19.38 0.718 4.58M 35.50G
+Table 2: Quantitative comparison on the real-world dehazing dataset RTTS. The best results are in bold.
 […]
-ENTROPY $\uparrow$ 7.164 6.907 7.080 7.090 7.079 7.112 7.143 7.405
-Performance on Homogeneous Haze.
-DehazeMatic achieves the SOTA in the average performance of indoor and outdoor scenes, and especially significantly outperforms the second-best method by 0.79 dB with fewer MACs in outdoor scenes.
-Figure 7 provides visual comparisons on the SOTS Indoor dataset [38].
-In the green and red boxed regions, DehazeMatic exhibits superior color consistency and effectively restores finer details, respectively.
-Trade-off between Performance and Overhead. As shown in Figure 8, our DehazeMatic achieves the best trade-off for both types of haze, demonstrating its high efficiency.
-Generalization on Real-world Haze Dataset.
-To evaluate generalization, we compare DehazeMatic with several competitive methods on the real-world haze dataset RTTS [38], using models trained on the RESIDE OTS dataset [38] for inference.
-As shown in Table 2 , our method outperforms all others across no-reference metrics, highlighting its superior generalization ability and practical applicability.
-4.3 Ablation Studies
+PSD (Chen et al. 2021) 0.920 27.713 4.598
+D4 (Yang et al. 2024b) 1.358 33.210 4.484
+DGUN (Mou, Wang, and Zhang 2022) 1.111 27.968 4.653
+RIDCP (Wu et al. 2023) 0.944 17.293 4.965
+CORUN (Fang et al. 2024a) 0.824 11.956 5.342
+SGDN (Fang et al. 2025) 0.873 11.549 5.128
+DehazeMatic 0.796 11.435 5.510
+Figure 5: Visual comparisons on homogeneous hazy images. The bottom displays zoomed-in error maps of selected regions, where darker blue indicates larger restoration errors. Please zoom in to view.
+Figure 6: Visual comparisons on non-homogeneous hazy images.
+The bottom displays zoomed-in error maps of selected regions, where darker blue indicates larger restoration errors.
+Please zoom in to view.
+Figure 7: Illustration of the trade-off between performance and runtime.
+The blue bar denotes runtime, while the green bar indicates performance.
+A larger gap, where the green bar exceeds the blue bar, signifies a better trade-off.
+Performance on Synthetic Homogeneous Haze
+As shown in Table 1, both DCMPNet (Zhang, Zhou, and Li 2024) and MB-TaylorFormer (Qiu et al.
+2023) are competitive approaches; however, our DehazeMatic achieves the best overall performance.
+DCMPNet utilizes depth information as guidance, while MB-TaylorFormer employs linear self-attention for feature extraction.
+These results demonstrate the superiority of jointly leveraging haze density maps and semantic maps for guidance, as well as simultaneously capturing both short- and long-range dependencies.
+A visual comparison is shown in Figure 5.
+Notably, only our method yields minimal restoration errors at object boundaries.
+Performance on Synthetic Non-homogeneous Haze
+As shown in Table 1, while MITNet (Shen et al.
+2023) and OKNet (Cui, Ren, and Knoll 2024) achieve competitive performance in PSNR and SSIM, respectively, our DehazeMatic consistently outperforms them by achieving SOTA results across all metrics.
+The error maps in Figure 6 demonstrate that our method preserves the most intricate details in the blue-boxed tree region and yields the smallest errors along the object boundaries in the red-boxed target area.
+Performance on Real-world Haze
+To evaluate the practicality and generalization ability of our model in real-world scenarios, we conduct experiments on the real-world haze dataset RTTS (Li et al.
+2018), comparing it with SOTA methods specifically designed for real-world dehazing.
+As shown in Table 2, our method outperforms all others in terms of no-reference metrics, demonstrating its superiority.
+Trade-off between Performance and Runtime
+To evaluate the efficiency of our model, we measure its runtime on an NVIDIA A100 GPU and compare the performance and runtime trade-off with the three most competitive methods, as shown in Figure 7.
+Although our method incorporates CLIP (Radford et al.
+2021), the additional inference time it introduces is negligible—only 3.8 ms.
+As a result, the overall runtime of our model is merely 30.03 ms, achieving the best trade-off.
+Moreover, our method reaches 33 frames per second (FPS), satisfying real-time processing requirements.
+Ablation Studies
 We conduct extensive ablation studies to demonstrate the effectiveness of each component.
@@
 Setting SOTS-Indoor NH-Haze
-$\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$
+[…]
 Only Short Dependencies 38.41 0.992 20.40 0.771
@@
 Conv. & Linear Attention 41.41 0.996 21.36 0.803
-Table 4: Ablation study of CLIP-enhanced Dual-path Aggregator.
+Explicit Coexistence of Short- and Long-range Dependencies
+To validate the effectiveness of the dual-stream design, we construct variant models by removing either the path responsible for capturing long-range dependencies or the one for short-range dependencies.
+The corresponding results are presented in the first three rows of Table 3.
+Furthermore, to validate the generality of our dual-stream design and show that it does not depend on specific feature extractors in each path, we construct a variant model that employs convolution (Zheng et al.
+2023) (i.e., short-range) and linear self-attention (Qiu et al.
+2023) (i.e., long-range), and repeat the above experiments.
+The corresponding quantitative results are reported in the last three rows of Table 3.
+These results confirm that the integration of complementary short- and long-range dependencies through the dual-stream design substantially improves image dehazing performance.
+Table 4: Ablation study of the CedA.
 Setting SOTS-Indoor NH-Haze
-$\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$
+[…]
 (a) Remove CedA Addition 39.55 0.993 20.71 0.781
@@
 DehazeMatic 41.50 0.996 21.47 0.806
-Explicit Coexistence of Short- and Long-range Dependencies.
-We remove either the Mamba path (i.e., long-range) or the Transformer path (i.e., short-range) in the TrambaDA blocks to construct variant models, while simultaneously excluding the CedA module.
-The corresponding results are shown in the first three rows of Table 3.
-Additionally, to verify that its superiority is not dependent on a specific feature extractor, we use convolution [78] (i.e., short-range), linear self-attention [51] (i.e., long-range) and the proposed CedA module to construct a variant model and repeat the above experiments.
-The corresponding quantitative results are presented in the last three rows of Table 3, demonstrating the versatility of our idea.
-Although linear self-attention [51] achieves better results than the adopted Mamba block, its time cost is significantly larger, so we do not adopt it.
-The comprehensive results in Table 3 demonstrate that the complementary short- and long-range dependencies introduced by the dual-path design, along with their reasonable aggregation, can significantly enhance image dehazing.
-Figure 9: Visual comparisons of haze density maps estimated using learned prompts and manually predefined prompts.
-CLIP-enhanced Dual-path Aggregator (CedA). As shown in Table 4, we conduct extensive ablation studies on the CedA to verify the effectiveness of each component.
+CLIP-enhanced Dual-path Aggregator (CedA)
+As shown in Table 4, we conduct extensive ablation studies on the CedA to verify the effectiveness of each component.
 (a) We replace the CedA with addition and concatenation operations to verify the necessity of adaptively aggregating short- and long-range dependencies in the dual paths.
-(b) We then verify the importance of jointly using haze density and semantic information for aggregation guidance by removing each component individually.
+(b) We then verify the importance of jointly using haze density and semantic information map for aggregation guidance by removing each component individually.
+Figure 8: Visual comparisons of haze density maps estimated by learned prompts and manually predefined prompts.
 (c) Next, we investigate the estimated haze density map.
-First, we verify the rationale for using the density map rather than other commonly used guidance signals in dehazing (i.e., the transmission map from DCP [22] and the depth map from Depth Anything [67]) to guide the aggregation.
+First, we verify the rationale for using the density map rather than other commonly used guidance signals in dehazing (i.e., the transmission map from DCP (He, Sun, and Tang 2010) and the depth map from Depth Anything (Yang et al.
+2024a)) to guide the aggregation.
 The results indicate that haze density is a more suitable influencing factor.
@@
 (d) Finally, we validate the necessity of each type of information in the refined semantic map by ablating either high-level semantic information (i.e., from CLIP) or low-level semantic information (i.e., input of Tramba blocks).
-5 Conclusion
-In this paper, we propose DehazeMatic, a framework that employs a dual-path design to explicitly and simultaneously capture complementary local fine-grained details and global contextual cues.
-Based on key insights from our experiments—namely that different tokens in an image require varying ranges of dependencies and that haze density and semantic information are the primary factors influencing the required range—we introduce the CLIP-enhanced Dual-path Aggregator.
-It leverages the rich semantic priors and powerful generalization ability encapsulated in CLIP to ensure that each dependency contributes optimally to performance while further exploring CLIP’s potential in image dehazing.
-Extensive experiments demonstrate that DehazeMatic outperforms SOTA methods on various benchmarks and holds promise for inspiring future research.
+Conclusion
+We propose DehazeMatic, demonstrating that explicitly capturing complementary short- and long-range dependencies, along with effective aggregation, can achieve SOTA performance and holds potential to inspire future research.
+To ensure that dependencies across different ranges contribute optimally, we first identify two key factors influencing their relative importance: haze density and semantic information.
+Guided by this insight, we introduce the CLIP-enhanced Dual-path Aggregator, which generates both maps to instruct the aggregation process.
+To the best of our knowledge, we are the first to accurately estimate fine-grained haze density maps, further unlocking CLIP’s potential in dehazing.